At what temperature will a 2.2 g sample of neon gas exert a pressure of 488 torr in a 3.7 L container?
Answer #1

Molar mass of Ne = 20.18 g/mol

mass(Ne)= 2.2 g

use:

number of mol of Ne,

n = mass of Ne/molar mass of Ne

=(2.2 g)/(20.18 g/mol)

= 0.109 mol

Given:

P = 488.0 torr

= (488.0/760) atm

= 0.6421 atm

V = 3.7 L

n = 0.109 mol

use:

P * V = n*R*T

0.6421 atm * 3.7 L = 0.109 mol* 0.08206 atm.L/mol.K * T

T = 266 K

Answer: 266 K